The Rec Centre has analysed its records for customer use of the squash courts and supercircuits classes. They found that the probability that a customer takes supercircuits classes is 0.20, the probability that a customer uses the squash courts given that they take supercircuit classes is 0.60, and the probability that if a customer uses the squash courts they also take supercircuits classes is 0.30.
The events of interest are defined as
SC for the event "customer takes supercircuits classes"
SCc for the event "customer does not take supercircuits classes"
SQ for the event "customer uses the squash courts"
SQc for the event "customer does not use the squash courts"
Match the words to the correct representation in probability symbols:
The probability that a customer uses the squash courts given that they take supercircuit classes is represented as
Answer 1
The probability that if a customer uses the squash courts they also take supercircuits classes is represented as
Answer 2
The probability that a customer takes supercircuits classes is represented as
Answer 3
Thank you so much

Answer by math_helper(2461)

These three items are given in the 2nd sentence:
P(SC) = 0.20
P(SQ | SC) = 0.60
P(SC | SQ) = 0.30
I am interpreting "if a customer uses the squash courts they also take supercircuit classes" as "given the customer uses the squash courts, they also take supercircuit classes" for if I interpret this as SC & SQ, the math does not work ( 0.30 does not equal (0.60)*(0.20) ).

1. P(SQ | SC) = 0.60
2. P(SC & SQ) = P(SQ | SC) * P(SC) = 0.60 * 0.20 = 0.12
3. P(SC) = 0.20

That you posted this question with the answers so readily available tells me you very likely need to increase your effort here.